OpenStudy (anonymous):

the right triangle is 20 sq cm , if one of the side is 5 cm.. what is the perimeter

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

If the perimeter of the square is 20 cm's then that means that each side of the square is 5 cm's. For the diagonal it cuts the square into 2 equal triangles so you can apply pythagorean's theory. a^2 + b^2 = c^2 a and b are the two sides of the square being 5 cm's and c is the diagonal. 5^2 + 5^2 = c^2 25 + 25 = c^2 50 = c^2 therefore the diagonal is the square root of 50 which is approx. 7

OpenStudy (solomonzelman):